OpenStudy (anonymous):

Show that square root of i= 1+i/square root of 2. Plz Help me as fast as u can!

OpenStudy (anonymous):

try to find what \[(\frac{1+ i }{ \sqrt{2} })^{2}\] is

OpenStudy (anonymous):

um go to math .com

OpenStudy (anonymous):

Do you know what i stands for?

OpenStudy (anonymous):

i = square root of -1

OpenStudy (anonymous):

Plz can you solve for me? I am in an urgent condition.

OpenStudy (anonymous):

Im having a hard time solving, I will send someone to help.

OpenStudy (anonymous):

@phi maybe able to help.

OpenStudy (phi):

Show that square root of i= 1+i/square root of 2. \[ \sqrt{i} = \frac{1+i}{\sqrt{2}} \] square both sides \[ i = \left(\frac{1+i}{\sqrt{2}} \right)^2 \] square the right-hand side. Treat i as an ordinary variable and this is just like the multiplication of the binomial (1+x)*(1+x). After doing the multiplication, replace any i^2 term with -1 and simplify